A ball falling freely in the field of gravity passed the final section of its trajectory of length 45 m during time interval 1 sec. Find the
height from which the ball fell if its initial velocity was zero. Give your answer in meters m]. Assume the free fall acceleration to be
10 m/s2. The air resistance is negligible.
1
Expert's answer
2020-11-03T10:28:29-0500

The distance trevelled by the free falling ball is:


"d_{last} = v_0t + \dfrac{gt_{last}^2}{2}"

where "v_0" is the initial speed of ball at the beginning of the last section, "g = 10m\/s^2" is the free fall acceleration, "t_{last} = 1s" is the time of covering the last section, "d_{last}= 45m" is the length of the last section.

Expressing "v_0", obtain:


"v_0 = d_{last} - \dfrac{gt_{last}^2}{2} = 45-\dfrac{10\cdot 1^2}{2} = 40m/s"

In order to gain such velocity from the rest, the ball should have spent the following amount of time:


"t = \dfrac{v_0}{g} = \dfrac{40}{10} = 4s"

In this time it covered the following distance:


"d = \dfrac{gt^2}{2} = \dfrac{10\cdot 4^2}{2} = 80m"

Thus, the total height from which the ball fell is:


"h = d + d_{last} = 80m + 45m = 125m"

Answer. 125 m.